### v1

# GShield: Mitigating Poisoning Attacks in Federated Learning

Abstract
Federated Learning (FL) has recently emerged as a revolutionary approach to collaborative training Machine Learning models. In particular, it enables decentralized model training while preserving data privacy, but its distributed nature makes it highly vulnerable to a severe attack known as Data Poisoning. In such scenarios, malicious clients inject manipulated data into the training process, thereby degrading global model performance or causing targeted misclassification. In this paper, we present a novel defense mechanism called GShield, designed to detect and mitigate malicious and low-quality updates, especially under non-independent and identically distributed (non-IID) data scenarios. GShield operates by learning the distribution of benign gradients through clustering and Gaussian modeling during an initial round, enabling it to establish a reliable baseline of trusted client behavior. With this benign profile, GShield selectively aggregates only those updates that align with the expected gradient patterns, effectively isolating adversarial clients and preserving the integrity of the global model. An extensive experimental campaign demonstrates that our proposed defense significantly improves model robustness compared to the state-of-the-art methods while maintaining a high accuracy of performance across both tabular and image datasets. Furthermore, GShield improves the accuracy of the targeted class by $43\%$ to $65\%$ after detecting malicious and low-quality clients.
December 2025
Keywords Federated Learning, Poisoning Attack, Privacy, Security.
1 Introduction
Federated Learning (FL, hereafter) is a new paradigm that enables the collaborative training of Machine Learning models across multiple decentralized devices or organizations while ensuring that raw data remains local and private [mcmahan2017communication]. By transmitting only model updates rather than sensitive data, Federated Learning aims to preserve user privacy, reduce communication costs, and support compliance with data protection regulations. At the same time, it leverages the diversity and richness of distributed data sources to build models that are more robust, accurate, and applicable across a wide range of real-world applications. While this design significantly enhances data privacy, it also raises new security vulnerabilities that are less severe in centralized learning. One of the most critical threats in FL is the poisoning attack [tolpegin2020data, sameera2024privacy]. In this kind of attack, malicious or compromised clients deliberately manipulate their local training data or model updates to corrupt the global learning process. Due to the distributed and often open nature of FL systems, the server typically lacks direct visibility into client data or training procedures, making it difficult to verify the integrity of individual updates. As a result, poisoned updates can be aggregated into the global model, leading to degraded performance or malicious behavior.
The challenge is particularly acute in realistic non-IID settings, where the natural heterogeneity of client data distributions causes benign updates to exhibit high variance. This variability significantly reduces the effectiveness of existing defense mechanisms, which often rely on simple distance-based outlier detection or robust aggregation rules that assume relatively homogeneous client behavior. As a result, poisoned updates can closely mimic legitimate ones, evade detection, and be aggregated into the global model, leading to degraded performance or targeted misclassification. These limitations motivate the need for a defense strategy that can accurately characterize benign client behavior under non-IID conditions and selectively filter malicious or low-quality contributions.
To address this challenge, we propose GShield, a defense mechanism that learns the distribution of benign client gradients through clustering and Gaussian modeling during an initial trusted phase. By establishing a reliable baseline of expected gradient patterns, GShield enables the server to identify and isolate anomalous updates while preserving the contributions of honest clients, thereby maintaining both robustness and high model accuracy in adversarial federated environments.
GShield operates in two distinct phases. During an initial Safe Round phase, the system assumes the absence of adversarial clients and collects gradient updates from participating clients. By applying cosine-similarity-based clustering, GShield identifies the dominant, high-density cluster of updates, which is presumed to represent benign behavior. Statistical characteristics of this cluster, specifically the mean and variance of similarity scores, are aggregated over multiple rounds to construct a Gaussian model that serves as a baseline representation of trusted client updates. This benign profile captures natural variability while filtering out outliers and low-quality contributions.
In the subsequent Anomaly Detection phase, GShield evaluates incoming client updates by comparing their gradient similarity scores against the learned benign distribution. Updates that conform to this distribution are selectively aggregated, while those exhibiting significant deviation are excluded from the global model update. This selective aggregation mechanism enables GShield to effectively isolate malicious or unreliable clients, preventing poisoned updates from influencing the global model. As a result, GShield preserves both model accuracy and robustness, even in the presence of targeted label-flipping attacks and severe data heterogeneity.
In our experimental evaluation, we compare the proposed GShield approach against several well-established and widely adopted defenses, including FLAME [nguyen2022flame], Krum [blanchard2017machine], Median [yin2018byzantine], and Trimmed Mean (TMean) [wang2025federated]. These methods are commonly used as baselines in the Federated Learning literature due to their demonstrated effectiveness across a wide range of attack models and threat scenarios [thein2024personalized, chelli2023fedguard, yang2025enhanced]. In line with this established evaluation practice, we include these representative techniques to ensure a fair, comprehensive, and consistent assessment of GShield’s robustness and performance.
By grounding anomaly detection in a statistically learned benign baseline rather than rigid heuristics or global assumptions, GShield provides a principled and adaptive defense against targeted poisoning attacks in federated learning, ensuring reliable model convergence and improved resilience across diverse datasets and attack scenarios.
In summary, the main contributions of this paper are as follows:
  • We propose GShield, a novel server-side defense for Federated Learning that mitigates targeted data poisoning attacks without requiring prior knowledge of adversaries, clean auxiliary data, or assumptions on client data distributions.
  • We introduce a statistically grounded approach that models benign client behavior through cosine-similarity-based clustering and Gaussian distribution learning, enabling reliable detection of malicious and low-quality updates under non-IID data.
  • We validate GShield through comprehensive experiments on tabular and image datasets, demonstrating superior robustness and significant improvements in targeted attack mitigation compared to state-of-the-art defenses.
The outline of this paper is as follows. In Section 2, we discuss the related papers in the state of the art. In Section 3, we describe the concepts useful to understand our approach. Section 4 presents our approach defense altogether the threat model. In Section 5, we present the experiments carried out to test our defense. Finally, Section 6 examines intriguing leads as future work and draws our conclusions.
2 Related Works
Recently, robust aggregation methods have been employed during the server-side aggregation phase to enhance resilience against Poisoning Attacks. Traditional approaches like FedAvg [mcmahan2017communication], which aggregate client updates using a simple arithmetic mean, have been shown to be vulnerable to such attacks. To address this, various robust aggregation techniques have been proposed to identify and mitigate the influence of malicious updates. For instance, the Median method replaces the mean with the element-wise median of client updates [yin2018byzantine], while the Trimmed Mean method removes a fixed proportion of the smallest and largest values before averaging [wang2025federated]. Norm Clipping constrains the impact of individual clients by bounding the norm of their updates to a predefined threshold [sun2019really]. The Krum algorithm selects a single client update whose distance to the majority of other updates is minimal, aiming to reflect consensus [blanchard2017machine]. Multi-Krum extends this by selecting and averaging the top-k closest updates, further enhancing robustness against outliers and adversarial contributions [blanchard2017machine]. In their research, the authors of [guerraoui2018hidden] introduced Bulyan, a robust aggregation method that combines MultiKrum and trimmed-mean techniques to defend against poisoning attacks. However, a key limitation of Bulyan is its assumption of IID (independent and identically distributed) data, which does not hold in many federated learning scenarios. To address malicious behavior, Fung et al. proposed FoolsGold [fung2020mitigating], which filters out adversarial updates by analyzing the cosine similarity of gradients, assuming that malicious clients tend to produce similar updates. In [pillutla2022robust], the authors presented Robust Federated Aggregation (RFA), which aggregates model updates using a smoothed Weiszfeld algorithm to compute a weighted geometric median, thereby preserving client data privacy while enhancing robustness. Other notable approaches include FLAME [nguyen2022flame], ZeKoC [chen2020zero], FLTrust [cao2022fltrust], ShieldFL [ma2022shieldfl], Adaptive Model Averaging [munoz2019byzantine], Residual-based Reweighting [fu2021attack], SEAR [zhao2021sear], and FedGuard [chelli2023fedguard], each offering distinct mechanisms to mitigate the impact of adversarial or unreliable updates.
In our experimental campaign, we compared our GShield approach with FLAME [nguyen2022flame], Krum [blanchard2017machine], Median [yin2018byzantine], and TMean [wang2025federated] methods. Prior works typically compare their proposed approaches with these techniques due to their effectiveness and popularity across various attack scenarios [thein2024personalized, chelli2023fedguard, yang2025enhanced]. Following this convention, we also include these representative methods in our experimental comparison to provide a fair and consistent evaluation of our proposed GShield.
3 Background
In this section, we introduce several foundational concepts that support our approach. Specifically, we explore the main characteristics and categories of Federated Learning (FL) and examine the main features of poisoning attacks in this context.
Table 1 summarizes the acronyms used in this paper.
Table 1: Summary of the acronyms used in the paper
Symbol Description
BA Backdoor Attack
DNN Deep Neural Network
DL Deep Learning
FL Federated Learning
GM Global Model
LFA Label Flipping Attack
LM Local Model
MIA Membership Inference Attack
ML Machine Learning
PM Poisoned Model
3.1 Federated Learning
Federated Learning (FL, hereafter) is a distributed approach developed to address optimization problems in an environment where a central server coordinates the learning process across a large number of decentralized devices. The devices, known as clients or workers, collaboratively train a Global Model (GM) by processing their local data samples and actively preserving data privacy by avoiding the exchange of raw data. We assume that each client $P_{i}$, for all $i\in[1,N]$, possesses a local training dataset $\mathcal{D}_{i}$. These datasets are mutually disjoint and follow a non-identically independent distributed (non-IID) pattern, i.e., $\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset$ for all $i,j\in[1,N]$ such that $P_{i}\neq P_{j}$. The FL objective seeks to minimize a global loss function $f(w)$, defined as the average of local objectives across $N$ clients:
$\min_{w}f(w)=\frac{1}{N}\sum_{i=1}^{N}F_{i}(w)$ (1)
Here, $w$ denotes the parameters of the shared global model. For simplicity, we represent both the input data and the model parameters as vectors. The local objective $F_{i}(w)$ for client $i$ typically corresponds to the empirical risk minimization, defined over its local data distribution $\mathcal{D}_{i}$ as:
$F_{i}(w)=\mathbb{E}_{x_{i}\sim\mathcal{D}_{i}}\left[\ell_{i}(w;x_{i})\right]$ (2)
where $\ell_{i}()$ is the local loss function and $x_{i}$ denotes the training samples. In this work, we consider $F_{i}(w)$ to be a non-convex function, which is optimized locally by each client using stochastic solvers such as Stochastic Gradient Descent (SGD). During each communication round, clients receive the current global model $w^{t}$ from the Aggregator Server, and perform $E$ epochs of local updates as follows:
$w_{i,E}^{t}=w_{i,0}^{t}-\eta\sum_{e=0}^{E-1}\nabla F_{i}\left(w_{i,e}^{t};\mathcal{B}_{i,e}\right)$ (3)
where $\eta$ is the local learning rate, and $\mathcal{B}_{i,e}$ denotes the mini-batch sampled at epoch $e$ on client $i$. The updated model $w_{i}^{t,E}$ is then sent to the central server for aggregation.
Figure 1 shows a general architecture of FL.
Figure 1: A general Federated Learning architecture
3.2 Poisoning Attacks on Federated Learning
Poisoning attacks aim to subvert the learning process by introducing malicious perturbations, either through crafted data samples or manipulated local model updates. Their objective is to degrade the performance or integrity of the global model. In Federated Learning (FL), clients collaboratively train a shared model by periodically transmitting local updates to a central server, which aggregates them to refine the global model. Due to the decentralized and trust-assumed nature of FL, a compromised client can inject adversarial updates into the aggregation process. These poisoned contributions propagate through the model, leading to convergence instability, biased predictions, or a significant drop in overall model accuracy. There exist two main categories of Poisoning Attacks, each exploiting different aspects of FL’s decentralized architecture, and requiring specific defense strategies, namely:
  • Data Poisoning Attacks, in which the training data used by local clients is targeted.
  • Model Poisoning Attacks, in which the local model updates sent to the server are manipulated.
Figure 2 illustrates the scheme of the two categories of Poisoning Attack in FL. During the FL training process, an attacker can execute a Poisoning Attack by modifying the client’s local data or directly modifying the local model parameters.
Figure 2: Example of Poisoning Attacks in FL
Beyond the specific techniques used, both Model and Data Poisoning Attacks can be further categorized based on the attacker’s intent. Broadly, these intents fall into two categories depending on the scope and specificity of the adversary’s objective: (i) Untargeted Poisoning, which aims at degrading the overall performance of the global model across all tasks, or (ii) Targeted Poisoning, which selectively impairing the model’s behavior on particular inputs or tasks. Additionally, there exists also a third category known as Semi-targeted Poisoning in which the adversary designates a specific source class (denoted as $C_{S}$) and aims to corrupt the GM such that instances from this class are misclassified as a different, attacker-chosen target class ($C^{*}$). The general formulation is given by:
$\arg\max_{m}f(x)_{m}={\left\{\begin{array}[]{ll}C^{*}&\text{if }y=C_{S}\\ y&\text{otherwise}\end{array}\right.}$
3.2.1 Data Poisoning Attacks
These attacks target the training data used by local clients, contaminating the local dataset during training. The can be categorized in the following types: (i) Label Flipping, (ii) Data Injection, and (iii) Backdoor techniques.
In Label Flipping Attack (LFA), the attacker flips labels of training samples (e.g., labeling a “fish” as a “dog”). Label-flipping attacks can be classified as clean-label and dirty-label variants. In clean-label scenarios, the adversary is constrained from modifying class labels, typically due to a certification mechanism that ensures label integrity and limits perceptible input manipulation. This attack can be perpetrated by modifying the samples in the training set and, for example, adding noise to the training data. Conversely, dirty-label attacks allow the adversary to inject training samples with intentionally incorrect labels—distinct from their true class—thereby inducing misclassifications toward a specific target class. In Data injection Attacks, a malicious actor tries to modify specific data samples within a dataset, injecting anomalous or irrelevant data points to skew local model updates. Finally, in Backdoor Attack involves an adversary embedding a covert mechanism into the model during training, enabling controlled manipulation of its predictions. This is achieved by associating specific, often imperceptible input patterns—known as triggers—with a target misclassification. The attack introduces a malicious perturbation $\Delta$ that activates only when the trigger is present, causing the model to misclassify the input while preserving standard behavior on clean, trigger-free data [liu2022backdoor, nicolazzo2025secure].
$F(X)=y_{i},\quad\forall X\sim\mathcal{D}_{\text{clean}}$ (4)
$F(X+\Delta)=y_{t},\quad\forall(X+\Delta)\sim\mathcal{D}_{\text{backdoor}}$ (5)
where $X$ represents the clean input data; $\Delta$ is the trigger pattern introduced by the attacker; $y_{i}$ denotes the true label of the clean input; and $y_{t}$ represents the attacker’s target label. The clean data is drawn from the distribution $\mathcal{D}_{\text{clean}}$, whereas the manipulated (backdoored) data comes from the distribution $\mathcal{D}_{\text{backdoor}}$.
3.2.2 Model Poisoning Attacks
These kinds of attacks manipulate the local model updates sent to the aggregator server. We can organize this type of attack into three categories [sameera2025federated], depending on how the adversary creates the models, namely: (i) Model Replacement, (ii) Optimization-based method, and (iii) Training Rule Manipulation.
In Model Replacement Attack, the attacker substitutes the models or gradients with crafted ones, achieved through random weights generation or gradient manipulation. In the second category of Optimization-based method, the adversary subtly alters the local training process (for instance, using different learning rates or objectives) to distort updates without obvious signs. Finally, in Training Rule Manipulation, the training rule is altered to introduce poison into the global model.
4 Overview of GShield
This section presents an overview of the proposed robust defense framework called GShield, designed to mitigate data poisoning attacks in the FL framework. We begin by examining the threat model related to targeted LFA in FL. Next, we introduce the GShield framework, highlighting its main components and the algorithms it employs. This is followed by a theoretical analysis of its effectiveness in mitigating poisoning attacks.
4.1 Threat Model
Targeted Label Flipping Attacks in FL are a type of Poisoning Attack where malicious clients attempt to corrupt the Global Model (GM) in such a way that specific inputs are misclassified into a target class chosen by the attacker. In particular, the adversary’s goal is to manipulate the local training data updates to force the GM to misclassify specific inputs (from the source class) as belonging to a specific target label, without significantly degrading overall model performance. This is different from untargeted attacks, which aim to reduce the overall accuracy or cause arbitrary misclassifications. We assume the Aggregator Server to be honest and non-compromised, and the attacker (that can be more than one entity) to have no control over the the Aggregator or the honest peers. In particular, as visible in Figure 3, the attackers poison their local training data by flipping the labels of training examples of a source class $C_{S}$ to a target class $C_{T}\in C$ while keeping the input data features unchanged. In brief, the steps of the attack are the following:
  • Every instance with the source label $C_{S}$ is changed to the target label $C_{T}$, to introduce intentional label noise into the training data.
  • The client trains the local model using this poisoned dataset for a fixed number of epochs.
  • The poisoned model update is sent to the Aggregator server during the standard federated aggregation step without being detected by the aggregation methods.
  • Over multiple rounds, the GM starts to misclassify inputs from the source class as the target class.
Figure 3: Our threat scenario
4.2 Defense Approach Description
This section presents an overview of the proposed robust defense framework called GShield, designed to mitigate label-flipping poisoning attacks in the FL system framework. Our proposed GShield functions without prior knowledge of attacker proportions, auxiliary datasets, or assumptions about the data distribution. It is deployed on the server side, aiming primarily to identify malicious clients and low-quality models present in the heterogeneous environment.
Algorithm 1 GShield: Defense against targeted poisoning attack
0: Global Model $\omega^{0}$, dataset $D_{i}$ for each client $i$, list of compromised clients AdvAdv
0: Final Global Model $\omega^{T}$ after $T$ rounds
1:for $t=1$ to $T$ do
2:  $S^{t}\leftarrow$ Randomly select $K$ clients
3:  Server sends $w^{t-1}$ to all clients in $S^{t}$
4:  for all client $i\in S^{t}$ in parallel do
5:   if $i\in Adv$ then
6:    Apply label flipping attack on $D_{i}$
7:   end if
8:   $g_{i}^{t}\leftarrow$ Compute gradients using SGD on local data $\nabla\mathcal{L}(w_{i}^{t-1};D_{i})$]
9:   Client $i$ sends $g_{i}^{t}$ to the server
10:  end for/*Detect benign clients using gradients*/
11:  $\mathcal{B}^{t}\leftarrow$ BenignClientSelection($\{g_{i}^{t}\}_{i\in S^{t}}$)
12:  $w^{t}\leftarrow$ Aggregate model from benign clients in $\mathcal{B}^{t}$
13:end for
14:return $\omega^{T}$
GShield is a server-side defense mechanism designed to detect and mitigate the influence of malicious and low-quality model updates in federated learning. It consists of two main phases:
  • Safe Phase. In the initial training rounds belonging to this phase GShield aims to learn and model benign client behavior by building a statistical baseline of trustworthy gradient updates.
  • Anomaly Detection Phase. During this phase, the system aims to detect and filter malicious or low-quality updates by comparing them against the learned benign baseline.
GShield begins with an initial phase, termed the Safe Round, during which it operates under the assumption that no adversarial clients are present. In this phase, GShield collects gradient updates from participating clients and applies clustering techniques to partition them into distinct groups. The dominant group, identified as the positive cluster, is presumed to comprise benign clients. In contrast, the remaining group, containing outliers and low-quality updates, is designated as the negative cluster. GShield then computes statistical measures, including the mean and standard deviation of the updates in the positive cluster, to construct a Gaussian Mixture Model (GMM). The global model is updated exclusively by aggregating the contributions from the positive cluster. This process is repeated for each round of the Safe Round phase, establishing a robust baseline that facilitates accurate anomaly detection and mitigation in subsequent training stages.
In the subsequent rounds, known as the Anomaly Detection Phase, GShield uses the previously learned GMM to assess incoming client updates. It filters out suspicious or inconsistent gradients by comparing them to the benign distribution established during the Safe Round phase. Only updates that conform to the expected benign behavior are aggregated. This selective aggregation process enables GShield to effectively identify and mitigate the influence of malicious or low-quality clients, thus preserving the integrity and performance of the global model throughout the training process. Figure 4 illustrates the overall workflow of GShield, particularly highlighting its operations during the Anomaly Detection Phase.
Figure 4: Proposed GShield receives local updates from clients and generates a new global model that eliminates poisoned local updates.
Algorithm 2 BenignClientSelection
0: Local Model $\omega^{0}$, current round $t$, safe round $T_{\text{safe}}$, threshold $z_{\alpha}$
0: Benign model $\mathcal{B}^{t}$
1:for all client $i\in S^{t}$ do
2:  $g_{i}^{t,L}\leftarrow$ Extract last-layer gradient
3:  Stack gradients $\{g_{i}^{t,L}\}$ into matrix $G^{t}$
4:end for
[…]
8:if $t\leq T_{\text{safe}}$ then
9:  $\mathcal{B}^{t}\leftarrow$ clients in $\mathcal{C}_{\text{max}}$
10:  $\mu_{safe},\sigma_{safe}\leftarrow$ mean $\mu_{t}$ and standard deviation of $Cos^{t}$ of $\mathcal{B}^{t}$
11:  if $t=T_{\text{safe}}$ then
12:   $\mu,\sigma\leftarrow$ Average of $\mu_{safe},\sigma_{safe}$
13:   Fit Gaussian model $\mathcal{N}(\mu,\sigma^{2})$
14:  end if
15:else
16:  $\mathcal{B}^{t}\leftarrow\emptyset$
17:  for all client $i\in S^{t}$ do
18:   Compute cosine similarity score $\text{sim}_{i}$
19:   Compute Mahalanobis distance $d_{i}=\frac{|\text{sim}_{i}-\mu|}{\sigma}$
20:   if $d_{i}\leq z_{\alpha}$ then
21:    $\mathcal{B}^{t}\leftarrow\mathcal{B}^{t}\cup\{i\}$ {Mark client as benign}
22:   end if
23:  end for
24:end if
25:return $\mathcal{B}^{t}$
We outline our method in Algorithm 1. At the beginning of round $t$, the server randomly selects a set $S^{t}$ of $K$ clients and sends the current global model $w^{t-1}$ to all clients in $S^{t}$. Each selected client $i\in S^{t}$ then performs local training on its dataset $D_{i}$. If client $i$ is an adversary (denoted as $Adv$), it applies a label-flipping attack to its dataset $D_{i}$. After completing the local training, each client computes its gradient $g_{i}^{t}=\nabla\mathcal{L}(w^{t-1};D_{i})$ using stochastic gradient descent and returns this gradient to the server.
Upon receiving the $K$ local gradients, the server implements a benign client detection procedure to identify which clients are likely behaving reliably. The server then aggregates the updates from these trustworthy clients to create the new global model $w^{t}$, while isolating any suspected malicious or low-quality clients to maintain the integrity of the global model.
The proposed Benign Client Selection procedure, presented in Algorithm 2, operates in two distinct phases. In each round $t$, the server first collects the last-layer gradients $g_{i}^{t,L}$ from every participating client $i\in S^{t}$ and constructs a gradient matrix $G^{t}$. The server then computes the cosine similarity $Cos^{t}$ across these gradient vectors and applies $k$-means clustering with $k=2$ to partition the clients into two clusters. The largest, densest cluster, denoted as $\mathcal{C}_{\mathit{max}}$, is presumed to comprise predominantly benign clients.
During the Safe Round phase ($t\leq T_{\mathit{safe}}$), the server treats all clients in $\mathcal{C}_{\mathit{max}}$ as benign and records their average cosine similarity $\mu_{t}$ and standard deviation $\sigma_{t}$. Upon reaching $t=T_{\mathit{safe}}$, it aggregates these statistics across the safe rounds to obtain global estimates of the benign distribution, $\mu$ and $\sigma$, and fits a corresponding Gaussian model $\mathcal{N}(\mu,\sigma^{2})$.
In the Anomaly Detection Phase ($t>T_{\mathit{safe}}$), the server evaluates new client updates by computing their cosine similarity scores $\mathit{sim}_{i}$ and calculating the associated Mahalanobis distance.
$d_{i}=\frac{\lvert\mathit{sim}_{i}-\mu\rvert}{\sigma}.$ (6)
Clients with $d_{i}\leq z_{\alpha}$ are classified as benign and added to $\mathcal{B}^{t}$. This procedure allows the server to reliably filter out malicious or low-quality clients, thereby ensuring robust and secure federated learning throughout the training process.
5 Experiments
In this section, we evaluate the performance of GShield compared to other existing approaches to assess its effectiveness and robustness. Specifically, we detail the experiment setup, including the employed datasets and the evaluation metrics, followed by a comprehensive presentation and critical analysis of the obtained results.
5.1 Datasets and Implementation Details
We extensively evaluated our proposed GShield method using publicly available benchmark datasets commonly adopted in this domain, encompassing both tabular and image-based modalities. For the tabular data, we considered two widely used network intrusion datasets: CIC-Darknet2020 and CSE-CIC-IDS2018 [hernandez2025intrusion]. To assess the generalizability of our approach beyond network traffic, we also included two image-based benchmarks: FashionMNIST and EMNIST-Digits. We configured FL with a total number of clients of $50$, where $50\%$ of the participants are considered for the aggregation. To evaluate the robustness and performance of the proposed GShield, we examined varying Poisoned Malicious Ratios (PMR), with $5\%$, $10\%$, $15\%$, $20\%$, and $25\%$ of the total clients configured as malicious participants. To control the degree of heterogeneity, we employed a Dirichlet distribution [hsu2019measuring] with a concentration parameter of $\alpha=0.5$ to generate non-IID training data for the clients.
System and machine setup. We conducted our experiments on an Ubuntu 22.04 LTS system equipped with an Intel Core i9 processor, 32 GB of RAM, and an NVIDIA Quadro P2000 GPU with 5 GB of GDDR5X memory. Our FL implementation was developed using the PyTorch framework, where clients were trained sequentially and independently in each global training round.
5.2 Evaluation Metrics
To assess the performance of our defense strategy, we mainly employed the following evaluation metrics:
  • Source Class Recall, which measures the proportion of correctly classified samples from the adversary’s chosen source class among all samples that truly belong to that class.
    $\text{SRecall}=\frac{1}{N_{\text{src}}}\sum_{i=1}^{N_{\text{src}}}(y_{i}=\hat{y}_{i})$
    where $N_{\text{src}}$ is the total number of instances belonging to the source class, $y_{i}$ is the true label of the $i$-th instance, and $\hat{y}_{i}$ is the predicted label.
  • F1-score, that is, the harmonic mean of precision and recall, ranging between 0 and 1.
    $\text{F1Score}=2\cdot\frac{\text{P}\cdot\text{R}}{\text{P}+\text{R}}$
    Precision $(P)$ indicates the proportion of correctly predicted positive samples among all samples predicted as positive, while Recall $(R)$ represents the proportion of correctly predicted positive samples among all actual positive samples.
5.3 Performance Analysis on Baseline Model
We conducted baseline evaluations under a non-poisoned scenario to analyze the F1 score of our FL configuration. Initially, the model was trained for $200$ communication rounds. From these results, we determined the optimal number of communication rounds for subsequent experiments. Figure 5 illustrates the performance of the global model in F1-score matrices across these communication rounds, with aggregation performed using the top-k participants.
We observed that the model started to converge after approximately $50$ communication rounds for both the CIC-Darknet2020 and CSE-CIC-IDS2018 datasets. Similarly, the model reached convergence in approximately $50$ rounds for both the FashionMNIST and EMNIST datasets. Consequently, we set the number of communication rounds to $100$ for tabular datasets and $50$ for image-based datasets in all subsequent experiments.
(a) CIC-Darknet2020
(b) CIC-IDS2018
(c) FashionMNIST
(d) EMNIST
Figure 5: Baseline.
Table 2: Comparison of proposed methods under different safe rounds on the tabular dataset. Bold indicates best results.
$SafeRound\downarrow$ $PMR\downarrow$ Krum TMean Median FLAME GShield
F1Score SRecall F1Score SRecall F1Score SRecall F1Score SRecall F1Score SRecall
CIC-Darknet2020
5 5% 0.7890 36.054 0.7144 14.582 0.7313 35.313 0.8366 36.213 0.8747 79.466
10% 0.8399 60.223 0.8373 76.016 0.8403 79.935 0.8448 74.751 0.8780 88.364
15% 0.8191 30.076 0.8244 27.157 0.7884 53.184 0.8031 44.598 0.8662 88.94
20% 0.7253 1.223 0.7162 1.094 0.7191 1.36 0.7328 1.166 0.8737 83.247
25% 0.7108 0.224 0.7012 0.201 0.7076 0 0.7207 0.216 0.8761 76.848
10 5% .8373 63.559 0.7428 18.518 0.7152 60.094 0.7915 37.891 0.8769 79.728
10% 0.8458 69.443 0.8399 75.737 0.8372 79.636 0.8392 74.229 0.8811 87.7
15% 0.8078 32.988 0.7997 31.724 0.8259 52.264 0.8220 45.472 0.8687 87.933
20% 0.7364 2.96 0.7201 1.937 0.7171 2.379 0.7291 2.542 0.8688 83.573
25% 0.7223 0.879 0.7075 0.394 0.7013 0.24 0.7085 0.72 0.8735 75.24
15 5% 0.8377 63.753 0.7592 69.41 0.7154 24.117 0.7956 39.895 0.8781 81.409
10% 0.8447 68.65 0.8164 40.176 0.8369 75.21 0.8391 74.092 0.8799 87.611
15% 0.8094 35.156 0.8113 38.155 0.8271 53.99 0.8237 48.013 0.8735 87.262
20% 0.7383 4.866 0.7237 3.687 0.7241 5.445 0.7327 4.157 0.8718 84.595
25% 0.7261 2.381 0.7115 1.595 0.6997 0.492 0.7119 1.792 0.8642 62.256
CSE-CIC-IDS2018
5 5% 0.974 86.314 0.9729 84.237 0.9714 67.756 0.9765 85.338 0.9821 91.094
10% 0.960 70.163 0.9676 71.057 0.9530 62.979 0.97147 72.567 0.9839 90.128
15% 0.965 53.57 0.9642 53.305 0.9466 41.851 0.9682 55.171 0.9822 86.265
20% 0.971 63.276 0.9601 51.724 0.9542 52.158 0.9683 54.125 0.9766 84.267
25% 0.975 53.999 0.9608 53.02 0.9562 59.868 0.9742 54.949 0.9805 83.716
10 5% 0.974 86.641 0.9728 84.273 0.9702 69.721 0.9780 85.648 0.9745 91.919
10% 0.960 72.561 0.9643 71.819 0.9512 74.435 0.9699 72.71 0.9773 89.754
15% 0.964 56.111 0.9615 52.925 0.9479 37.903 0.9663 55.677 0.9739 87.239
20% 0.931 54.547 0.9597 52.576 0.9542 56.327 0.9641 55.584 0.9829 84.032
25% 0.956 56.626 0.9692 52.155 0.9534 52.814 0.9706 54.114 0.9758 89.977
15 5% 0.966 86.321 0.9728 84.739 0.9713 71.413 0.9789 86.332 0.9800 91.276
10% 0.960 71.473 0.9634 70.516 0.9515 75.897 0.9691 77.834 0.9739 89.513
15% 0.949 62.773 0.9597 57.298 0.9476 55.617 0.9654 57.122 0.9817 86.213
20% 0.940 56.86 0.9577 53.013 0.9535 59.302 0.9617 54.712 0.9773 85.412
25% 0.948 57.709 0.9674 52.405 0.9529 50.01 0.9681 58.111 0.9740 91.213
Table 3: Comparison of the proposed method GShield under different safe rounds on the image dataset. Bold indicates best results.
$SafeRound\downarrow$ $PMR\downarrow$ Krum TMean Median FLAME GShield
F1Score SRecall F1Score SRecall F1Score SRecall F1Score SRecall F1Score SRecall
FashionMNIST
5 5% 0.854 94.456 0.831 93.306 0.865 94.958 0.870 95.554 0.878 95.922
10% 0.852 91.480 0.848 93.110 0.864 92.602 0.865 92.438 0.876 93.614
15% 0.852 91.030 0.852 93.846 0.867 93.158 0.871 92.274 0.872 93.226
20% 0.856 86.678 0.827 92.902 0.866 90.432 0.865 88.274 0.868 93.326
25% 0.837 80.516 0.840 88.750 0.847 83.500 0.851 84.984 0.863 92.280
10 5% 0.873 94.36 0.831 93.31 0.865 95.00 0.868 95.55 0.877 95.90
10% 0.872 91.42 0.848 92.11 0.865 92.61 0.864 92.46 0.874 93.71
15% 0.872 90.83 0.852 92.85 0.866 93.22 0.870 92.47 0.873 93.56
20% 0.865 85.93 0.827 90.90 0.867 91.12 0.866 89.89 0.871 92.60
25% 0.856 80.16 0.840 88.75 0.848 85.28 0.855 887.77 0.858 90.17
15 5% 0.875 95.64 0.833 97.24 0.866 95.35 0.870 96.06 0.878 97.27
10% 0.872 93.40 0.846 93.29 0.864 93.16 0.864 93.63 0.871 94.64
15% 0.874 94.07 0.851 94.17 0.868 94.36 0.869 94.28 0.872 95.19
20% 0.871 92.78 0.830 93.18 0.869 93.70 0.870 93.45 0.872 94.53
25% 0.867 91.52 0.840 90.02 0.856 92.54 0.859 93.71 0.866 93.98
EMNIST
5 5% 0.9808 95.22 0.9845 95.481 0.9601 95.711 0.9881 96.975 0.9851 99.193
10% 0.9774 94.382 0.9812 94.142 0.9403 93.688 0.9883 95.471 0.9842 98.891
15% 0.9745 88.793 0.9718 88.918 0.9152 86.424 0.9848 89.989 0.9834 99.064
20% 0.9727 88.802 0.9671 88.951 0.9016 88.218 0.9835 88.725 0.9824 98.343
25% 0.9760 86.77 0.9799 85.046 0.8988 85.731 0.9832 87.278 0.9839 97.949
10 5% 0.979 95.582 0.985 95.623 0.960 95.285 0.988 96.889 0.984 98.022
10% 0.978 94.077 0.981 93.325 0.943 92.239 0.988 95.391 0.987 95.93
15% 0.975 88.817 0.972 89.379 0.916 87.356 0.985 90.24 0.987 93.519
20% 0.974 86.638 0.963 88.334 0.905 87.895 0.983 88.731 0.990 91.57
25% 0.974 85.145 0.963 85.846 0.909 84.908 0.985 87.047 0.983 91.903
15 5% 0.979 95.796 0.976 95.231 0.974 95.374 0.988 96.894 0.984 97.27
10% 0.973 88.597 0.971 94.352 0.960 94.302 0.987 95.373 0.981 97.924
15% 0.978 86.866 0.948 89.585 0.882 89.838 0.984 90.322 0.980 93.743
20% 0.978 86.866 0.925 87.088 0.955 85.582 0.982 89.024 0.989 92.986
25% 0.978 86.866 0.979 84.849 0.954 84.968 0.984 86.741 0.983 88.391
5.4 Performance Analysis of GShield in Tabular Dataset
In this section, we examine the robustness of the proposed GShield method under various safe zones configurations and compare its performance with that of the state-of-the-art method. To assess the sensitivity to the safe period, we configured $SafeRound$ to 5, 10, 15. We used the same experimental setup for all state-of-the-art approaches. Specifically, no malicious participants were introduced until the prescribed $SafeRound$, after which malicious participants were added in subsequent rounds. Table 2 presents the performance of our GShield framework compared to state-of-the-art methods on two tabular benchmark datasets under varying PMR attack ratios and different safe rounds for brevity. GShield consistently outperforms other methods across all datasets and attack scenarios. For CIC-Darkent2020, at lower safe rounds (e.g., 5 rounds), all approaches exhibit relatively low recall values due to the difficulty of detecting malicious updates in a highly non-IID environment. Nevertheless, the proposed GShield demonstrates effectiveness, achieving a 5.4% higher recall than the best baseline (FLAME) at 10% PMR, highlighting its enhanced capability to identify adversarial clients even during early aggregation phases.
As the number of safe rounds increases (e.g., 10–15), the overall recall of all methods improves since additional aggregation rounds provide better global context for anomaly detection. Importantly, GShield maintains its robustness and detection stability, achieving $90.55\%$ recall at 15 safe rounds and $10\%$ PMR. While the numerical gain (approximately $1.6\%$) over the next best approach may appear modest, it reflects a significant consistency advantage, demonstrating that GShield sustains high detection accuracy even as the model converges and attack strength intensifies. This stable performance across varying conditions validates GShield’s resilience and its adaptability to heterogeneous, non-IID federated learning environments.
For the CSE-CIC-IDS2018 dataset, we observed that GShield achieved approximately $64.884\%$ to $43.253\%$ higher performance for the $SRecall$ than other methods when $SafeRound$ was set to 5 with a PMR of $5\%$. In the case of a $safeRound$ of 5 with a PMR of $25\%$, the $SRecall$ for GShield showed an improvement of approximately $76\%$ compared with other methods, while all other approaches were closer to $0\%$. This result demonstrates GShield’s ability to detect malicious participants effectively while maintaining the $SRecall$. Similarly, in this setting, GShield maintained an $F1Score$ of approximately $87\%$ across all cases, which was higher than the other approaches, whose scores ranged from $70\%$ to $72\%$. A similar trend was observed in other cases as well. The proposed GShield maintained the performance of the global model by effectively detecting both malicious and low-quality models during each round. Under the worst-case attack setting ($SafeRound=15$ and $PMR=15\%$), the $SRecall$ reached $91.213\%$, which is approximately $33\%$ to $41\%$ higher than that of other methods. Meanwhile, the global model achieved an $F1Score$ of $0.9740$.
5.5 Performance Analysis of GShield in Image Datasets
Table 3 illustrates the performance of the proposed GShield method under different $SafeRound$ and attack scenario settings. In the FashionMNIST dataset, for the $SRecall$ metric, GShield achieved approximately $2–3\%$ improvement over the competing methods for $PMR=5\%$ and $SafeRound=5$. Similarly, we observed that the $F1Score$ of the proposed GShield and FLAME remained almost identical. However, compared to other approaches, GShield improved approximately $2–4\%$. As the poisoning intensity increased, the $SRecall$ of all approaches declined significantly, although most methods attempted to maintain a stable $F1Score$. Notably, GShield still demonstrated an improvement of about $4–12\%$ when the $PMR$ increased to $25\%$, highlighting its robustness in handling higher levels of data poisoning.
Furthermore, as the $SafeRound$ increased, GShield consistently outperformed the other approaches. However, in most cases, FLAME’s performance was comparable to GShield, primarily because FLAME is capable of distinguishing anomalous models even under a high proportion of malicious participants. Nevertheless, GShield outperformed all other methods in both metrics, achieving approximately $1–3\%$ higher performance in $SRecall$, while maintaining a comparable $F1Score$ under the worst-case attack scenario, where other methods exhibited significant degradation.
5.6 Performance Analysis of GShield Effectiveness with Differential Privacy
To further investigate the effectiveness of GShield when integrated with Differential Privacy (DP), we compared its performance against several state-of-the-art defense frameworks under the same attack scenarios. In this experiment, $25\%$ of the participating clients were designated as malicious, each poisoning $50\%$ of their local samples, and the SafeRound was set to $10$ to evaluate the system’s resilience under these conditions. The objective was to assess how well GShield preserves model accuracy and robustness while maintaining privacy protection. As shown in Figure 6, GShield consistently outperformed existing methods in mitigating the impact of poisoning attacks, even under differential privacy constraints. This demonstrates that the incorporation of DP does not significantly compromise the model’s learning capability. Moreover, GShield achieved faster convergence and maintained stable performance across different attack scenarios, highlighting its ability to balance both privacy preservation and defense robustness more effectively than existing approaches.
(a) CIC-Darknet2020
(b) CIC-IDS2018
(c) FashionMNIST
Figure 6: Performance comparison demonstrating GShield’s robustness under Differential Privacy.
5.7 Computational Overhead of GShield
In this section, we discuss the computation overhead of the proposed approach against other state-of-the-art methods, whose results are visible in Table 4. The result illustrated that our proposed GShield has lower computation compared with Krum and FLAME. For the FashionMNIST dataset, GShield ($0.355$ seconds) is $80.9\%$ faster than FLAME ($1.859$ seconds) and $49.6\%$ faster than Krum ($0.705$ seconds). A similar trend is observed across other datasets. On CIC-Darknet2020, GShield reduces computation costs by $69.9\%$ compared to FLAME and by $67.7\%$compared to Krum. Additionally, on the EMNIST dataset, GShield shows a $78.1\%$ improvement over FLAME. The TMean and Median have the lowest computational cost, ranging from $0.002$ to $0.038$ seconds. However, their simplicity comes from basic statistical aggregation methods that do not address adversarial manipulation. As a result, these approaches can be easily compromised by targeted poisoning strategies. In contrast, GShield provides robustness against malicious updates with only a slight increase in runtime compared with this approach. This ensures effective security in FL systems by filtering out malicious model updates without imposing significant latency on the aggregation process.
Table 4: Comparison of GShield’s computational time (in seconds) with the state-of-the-art approach.
[…]
CIC-Darknet2020 CSE-CIC-IDS2018 FashionMNIST EMNIST
Krum [blanchard2017machine] 0.372 0.216 0.705 0.659
TMean [yin2018byzantine] 0.003 0.003 0.020 0.038
Median [yin2018byzantine] 0.002 0.004 0.019 0.031
FLAME [nguyen2022flame] 0.399 0.863 1.859 1.143
GShield (proposed) 0.120 0.138 0.355 0.250
6 Conclusion
In this paper, we introduced GShield, a robust server-side defense mechanism designed to mitigate targeted data poisoning attacks in Federated Learning, particularly under non-IID data distributions. By learning a statistical representation of benign client behavior through gradient clustering and Gaussian modeling, GShield effectively identifies and filters malicious and low-quality updates without requiring prior knowledge of adversaries or auxiliary datasets. Extensive experimental results demonstrate that GShield significantly improves model robustness and preserves high accuracy across both tabular and image datasets when compared to state-of-the-art defenses.
We plan to extend our work in the future to handle more adaptive and colluding adversaries, explore its effectiveness against other attack types such as model replacement and backdoor attacks, and investigate dynamic or continual updating of the benign profile to further enhance robustness in long-running Federated Learning deployments.

### v2

@@
 These datasets are mutually disjoint and follow a non-identically independent distributed (non-IID) pattern, i.e., $\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset$ for all $i,j\in[1,N]$ such that $P_{i}\neq P_{j}$.
-The FL objective seeks to minimize a global loss function $f(w)$, defined as the average of local objectives across $N$ clients:
+The FL objective seeks to minimize a global loss function $f(w)$, defined as the average of local objectives across NN clients:
 $\min_{w}f(w)=\frac{1}{N}\sum_{i=1}^{N}F_{i}(w)$ (1)
@@
 $w_{i,E}^{t}=w_{i,0}^{t}-\eta\sum_{e=0}^{E-1}\nabla F_{i}\left(w_{i,e}^{t};\mathcal{B}_{i,e}\right)$ (3)
-where $\eta$ is the local learning rate, and $\mathcal{B}_{i,e}$ denotes the mini-batch sampled at epoch $e$ on client $i$.
+where $\eta$ is the local learning rate, and $\mathcal{B}_{i,e}$ denotes the mini-batch sampled at epoch $e$ on client ii.
 The updated model $w_{i}^{t,E}$ is then sent to the central server for aggregation.
@@
 Each selected client $i\in S^{t}$ then performs local training on its dataset $D_{i}$.
-If client $i$ is an adversary (denoted as $Adv$), it applies a label-flipping attack to its dataset $D_{i}$.
+If client $i$ is an adversary (denoted as AdvAdv), it applies a label-flipping attack to its dataset $D_{i}$.
 After completing the local training, each client computes its gradient $g_{i}^{t}=\nabla\mathcal{L}(w^{t-1};D_{i})$ using stochastic gradient descent and returns this gradient to the server.
@@
 In this section, we examine the robustness of the proposed GShield method under various safe zones configurations and compare its performance with that of the state-of-the-art method.
-To assess the sensitivity to the safe period, we configured $SafeRound$ to 5, 10, 15.
+To assess the sensitivity to the safe period, we configured SafeRoundSafeRound to 5, 10, 15.
 We used the same experimental setup for all state-of-the-art approaches.
-Specifically, no malicious participants were introduced until the prescribed $SafeRound$, after which malicious participants were added in subsequent rounds.
+Specifically, no malicious participants were introduced until the prescribed SafeRoundSafeRound, after which malicious participants were added in subsequent rounds.
 Table 2 presents the performance of our GShield framework compared to state-of-the-art methods on two tabular benchmark datasets under varying PMR attack ratios and different safe rounds for brevity.
@@
 This stable performance across varying conditions validates GShield’s resilience and its adaptability to heterogeneous, non-IID federated learning environments.
-For the CSE-CIC-IDS2018 dataset, we observed that GShield achieved approximately $64.884\%$ to $43.253\%$ higher performance for the $SRecall$ than other methods when $SafeRound$ was set to 5 with a PMR of $5\%$.
-In the case of a $safeRound$ of 5 with a PMR of $25\%$, the $SRecall$ for GShield showed an improvement of approximately $76\%$ compared with other methods, while all other approaches were closer to $0\%$.
-This result demonstrates GShield’s ability to detect malicious participants effectively while maintaining the $SRecall$.
+For the CSE-CIC-IDS2018 dataset, we observed that GShield achieved approximately $64.884\%$ to $43.253\%$ higher performance for the SRecallSRecall than other methods when SafeRoundSafeRound was set to 5 with a PMR of $5\%$.
+In the case of a safeRoundsafeRound of 5 with a PMR of $25\%$, the SRecallSRecall for GShield showed an improvement of approximately $76\%$ compared with other methods, while all other approaches were closer to $0\%$.
+This result demonstrates GShield’s ability to detect malicious participants effectively while maintaining the SRecallSRecall.
 Similarly, in this setting, GShield maintained an $F1Score$ of approximately $87\%$ across all cases, which was higher than the other approaches, whose scores ranged from $70\%$ to $72\%$.
@@
 The proposed GShield maintained the performance of the global model by effectively detecting both malicious and low-quality models during each round.
-Under the worst-case attack setting ($SafeRound=15$ and $PMR=15\%$), the $SRecall$ reached $91.213\%$, which is approximately $33\%$ to $41\%$ higher than that of other methods.
+Under the worst-case attack setting ($SafeRound=15$ and $PMR=15\%$), the SRecallSRecall reached $91.213\%$, which is approximately $33\%$ to $41\%$ higher than that of other methods.
 Meanwhile, the global model achieved an $F1Score$ of $0.9740$.
 5.5 Performance Analysis of GShield in Image Datasets
-Table 3 illustrates the performance of the proposed GShield method under different $SafeRound$ and attack scenario settings.
-In the FashionMNIST dataset, for the $SRecall$ metric, GShield achieved approximately $2–3\%$ improvement over the competing methods for $PMR=5\%$ and $SafeRound=5$.
+Table 3 illustrates the performance of the proposed GShield method under different SafeRoundSafeRound and attack scenario settings.
+In the FashionMNIST dataset, for the SRecallSRecall metric, GShield achieved approximately $2–3\%$ improvement over the competing methods for $PMR=5\%$ and $SafeRound=5$.
 Similarly, we observed that the $F1Score$ of the proposed GShield and FLAME remained almost identical.
 However, compared to other approaches, GShield improved approximately $2–4\%$.
-As the poisoning intensity increased, the $SRecall$ of all approaches declined significantly, although most methods attempted to maintain a stable $F1Score$.
-Notably, GShield still demonstrated an improvement of about $4–12\%$ when the $PMR$ increased to $25\%$, highlighting its robustness in handling higher levels of data poisoning.
-Furthermore, as the $SafeRound$ increased, GShield consistently outperformed the other approaches.
+As the poisoning intensity increased, the SRecallSRecall of all approaches declined significantly, although most methods attempted to maintain a stable $F1Score$.
+Notably, GShield still demonstrated an improvement of about $4–12\%$ when the PMRPMR increased to $25\%$, highlighting its robustness in handling higher levels of data poisoning.
+Furthermore, as the SafeRoundSafeRound increased, GShield consistently outperformed the other approaches.
 However, in most cases, FLAME’s performance was comparable to GShield, primarily because FLAME is capable of distinguishing anomalous models even under a high proportion of malicious participants.
-Nevertheless, GShield outperformed all other methods in both metrics, achieving approximately $1–3\%$ higher performance in $SRecall$, while maintaining a comparable $F1Score$ under the worst-case attack scenario, where other methods exhibited significant degradation.
+Nevertheless, GShield outperformed all other methods in both metrics, achieving approximately $1–3\%$ higher performance in SRecallSRecall, while maintaining a comparable $F1Score$ under the worst-case attack scenario, where other methods exhibited significant degradation.
 5.6 Performance Analysis of GShield Effectiveness with Differential Privacy